I’m confused on how to do the question. Please help
Lecture 15 At-home Practice Buffer conceptual and calculation: A chemist needs a pH 3.90 buffer. Make the best choice from these four acids to start making a buffer nearest the desired pH. * Hydrofluoric acid, HF Ka-6.8x104 . Formic acid, НСНО, K,-1.7x10-4 Propionic acid, HC,HO2 K, 1.3x105 Acetic acid, HCH,CO2 K, 1.7x10 Assume the best acid is combined with its appropriate conjugate base. What is the [base]/[acid] ratio needed to make the pH 3.90 buffer? (Format the answer as base:1 acid.) . Titration calculation A chemist uses 45.36 mL of 1.500 M sodium hydroxide to titrate 50.00 mL of aqueous phosphoric acid solution. At the titration endpoint, the reaction product is sodium phosphate. Calculate the concentration of the analyte in the original solution.
Answer #1

A solution which contatins weak acid with its salt (conjugate base) with strong base and vice versa is called a buffer solution. We are given weak acids as their Ka value is small

Calculating pKa value from Ka value by pKa = -log(Ka) we have -

HF ----- -log(6.8*10-4) = 3.168

HCHO2 ------ -log(1.7*10-4) = 3.77

HC3H5O2 ------ -log(1.3*10-5) = 4.89

HCH3CO2 ------ -log(1.7*10-5) = 4.77

The buffer action is best if pKa-1 < pH < pKa+1

Also, the pH of a buffer solution is given by henderson hasselbalch equation -

pH = pKa + log(conjugate base/acid) where (conjugate base)/acid is the the ratio of concentration of (conjugate base) and acid.

We want a pH of 3.90 and the closest value of pKa is 3.77 that is of formic acid.

So, we will use that

so, 3.9 = 3.77 + log(base/acid) so, log(base/acid) = 0.13 or base/acid = 100.13 = 1.3489

so, it will be 1.3489 base : 1 acid